OpenStudy (anonymous):

Differentiate y=ln(x-1/X+1)

OpenStudy (anonymous):

y = ln(x/x-1) .........x=ln(y/y-1) e^x = y/(y-1) ........(y-1)e^x = y ye^x-e^x = y .........ye^x - y = e^x y(e^x -1) = e^x

OpenStudy (ksaimouli):